OpenStudy (anonymous):

Andy has $150. Marta has $120. In how many years will they both have the same amount of money if Andy saves $110 per year and Marta saves $120? Andy $150 $260 $370 $480 Marta$120 $240 $360 $480 Year 0 1 2 3 Here is a problem that can also be solved by making a table. Josie has $80. Addie has $30. In how many months will they both have the same amount of money if Josie saves $30 per month and Addie saves $40? PLEASE HELP!!!! Please Please Please!

OpenStudy (texaschic101):

150 + 110y = 120 + 120y 110y - 120y = 120 - 150 -10y = -30 y = 3 years check.. 150 + 110(3) = 120 + 120(3) 150 + 330 = 120 + 360 480 = 480 (correct) Andy and Marta will have the same amount of money in 3 years ================= 80 + 30m = 30 + 40m 30m - 40m = 30 - 80 -10m = - 50 m = 5 check... 80 + 30(5) = 30 + 40(5) 80 + 150 = 30 + 200 230 = 230 (correct) Josie and Addie will have the same amount of money in 5 months any questions ?

OpenStudy (anonymous):

Thanks So Much. I Dont Have Any Questions.

OpenStudy (texaschic101):

glad to help :)